Trey has 9 new magazines to read. Let M be the number of magazines he would have left to read after reading R of them. Write an equation relating M to R. Then graph your equation using the axes below.




1
Expert's answer
2022-02-21T17:00:11-0500

The equation is "M+R=9, 0\le R\le 9"